# Exploring Federated Unlearning: Analysis, Comparison, and Insights

### v4

@@ I Introduction
Federated learning is a distributed machine learning paradigm where clients’ devices collaborate to train a model without sharing their local data. Each client trains the model locally using its data and then sends model updates to a central server. The central server collects all received updates to train a global model and then sends the global model to clients for further training. This iterative process continues until the model achieves the desired performance. However, safeguarding federated learning clients from privacy attacks, such as membership inference, which could potentially leak private information, remains a pressing challenge. Furthermore, with the emergence of new data regulations such as the European Union’s General Data Protection Regulation (GDPR), clients are granted the “right to be forgotten”, enabling them to request the removal of their private data from models even after the data have been utilized for training a global model. A straightforward approach to accomplishing this is retraining the model from scratch without the removed data. However, this approach imposes a considerable computational burden. Moreover, when implementing unlearning in a decentralized manner, additional challenges arise. First, how to minimize the effect of unlearning on the accuracy of the federated learning model? Second, how can we guarantee that the client’s data is effectively erased as requested? Federated unlearning is crucial for removing obsolete data to keep models relevant and accurate in dynamic environments, such as personalized recommendations or autonomous systems.
@@
Federated unlearning uniquely addresses the challenge of selectively removing data from federated learning systems to comply with privacy regulations. Unlike traditional methods such as device anonymity, secure aggregation, differential privacy, and SMPC, which protect data during training but lack mechanisms for retrospective removal, federated unlearning enables precise data erasure while maintaining model integrity. As illustrated in Figure 2, federated unlearning requests typically fall into three levels: class-level, client-level, and sample-level. Specifically, class-level unlearning removes an entire class of data from a model, client-level unlearning removes all data associated with a particular client’s device, and sample-level unlearning focuses on removing specific data points’ impact from a model. Up to now, various federated unlearning methods have been proposed, and we analyze and compare these methods in this paper. We provide insights to assist in understanding their effectiveness from the following perspectives :
@@ II-B Applications
  • Telecommunications and Smart City. Federated unlearning is crucial in telecommunications and smart city applications, addressing scalability, real-time processing, and diverse data sources. Frameworks such as $\text{FedME}^{2}$ [9] and FedFilter [15] enhance secure communications by enabling selective erasure of outdated or incorrect data, optimizing system performance. Specifically, FedFilter uses edge caching to accelerate federated unlearning, minimizing latency for real-time Internet of Vehicles operations in smart cities, thereby enhancing resilience against security threats and privacy breaches [15].
@@ IV Experiments
TABLE II: Experimental results for client-level federated unlearning methods on MNIST, CIFAR-10 and SVHN datasets. [^2] GitHub for Codes.

### v5

# Exploring Federated Unlearning: Review, Comparison, and Insights

@@ I Introduction
Federated learning is a distributed paradigm in which clients independently train a model on their local datasets and subsequently transmit updates, rather than raw data, to a central server. The server aggregates these updates into a global model, which is then shared back to the participating clients for further iterations. Despite preserving data locally, this process can unintentionally disclose sensitive information through the gradient updates exchanged between clients and the server, thus enabling membership inference attacks. Such vulnerabilities compromise user privacy, and conflict with regulatory requirements, underscoring the need for robust safeguards in federated learning environments. Moreover, with the emergence of new data regulations such as the European Union’s General Data Protection Regulation (GDPR), clients are granted the “right to be forgotten”, enabling them to request the removal of their private data from models even after the data have been utilized for training a global model. A straightforward approach to accomplishing this is retraining the model from scratch without the removed data. However, this approach imposes a considerable computational burden. Additionally, when implementing unlearning in a decentralized manner, additional challenges arise. First, how to minimize the effect of unlearning on the accuracy of the federated learning model? Second, how can we guarantee that the client’s data is effectively erased as requested? Federated unlearning is crucial for removing obsolete data to keep models relevant and accurate in dynamic environments, such as personalized recommendations or autonomous systems.
@@
Federated unlearning addresses the need to selectively remove data from federated learning systems in order to comply with privacy regulations. Unlike conventional techniques (e.g., device anonymity, secure aggregation, differential privacy, and SMPC) that safeguard data during training but lack retrospective removal capabilities, federated unlearning enables fine-grained erasure while preserving model integrity. As illustrated in Figure 2, federated unlearning requests typically target three levels: class-level (removing all data within a class), client-level (removing all data originating from a specific client), and sample-level (removing the influence of particular data points). Notably, while the federated unlearning framework supports the removal of client data, it specifically focuses on mitigating the residual impact of such data on the learned model rather than physically deleting raw data from client devices. Numerous federated unlearning methods have been proposed, and we provide insights from the following perspectives :
@@ II-B Applications
  • Telecommunications and Smart City. Federated unlearning is crucial in telecommunications and smart city applications, addressing scalability, real-time processing, and the diversity of data sources. FedFilter [15] framework enhances secure communications by enabling the selective erasure of outdated or incorrect data, thereby optimizing system performance and trustworthiness. In particular, it uses edge caching to expedite federated unlearning by offloading model updates closer to the data, and applies user‐clustering techniques (e.g., sparsity‐based similarity) to deliver targeted personalization. Its hierarchical aggregation strategy shares only “base” model layers globally, keeping specialized “personalized” layers local to each cluster, which lowers communication overhead and mitigates the risk of data leakage. The unlearning component then removes malicious or obsolete inputs directly at or near the edge, curbing any negative impact on system accuracy and security. By localizing these processes, FedFilter reduces privacy risks and latency for real‐time IoV operations in smart cities, ensuring robust performance even in large‐scale, heterogeneous environments [15].
@@ IV Experiments
TABLE II: Experimental results for client-level federated unlearning methods on MNIST, CIFAR-10 and SVHN datasets. [^2] GitHub Repository.